Caan Corporation produces industrial robots for high-precision manufacturing. The following information is given for Caan Corporation:
Question 7 Caan Corporation produces industrial robots for high-precision manufacturing. The following information is given f(c) x Your answer is incorrect. Try again. Calculate the target selling price. 1589 Target selling price $L

I am not sure for part C,

Thank you.
Answer #1

rate positively ..

Target selling price = per unit fixed cost + per unit variable cost + per unit ROI required
i Asset invested =         51,145,000
ii ROI % = 25%
iii=i*ii ROI $=         12,786,250
iv Number of unit = 2650
v=iii/iv Profit required per unit                  4,825
vi Variable cost per unit 804
380+290+78+56
vii Fixed cost per unit = 737
620+117
viii=vi+vii+v Target selling price =                  6,366
Ans =                  6,366